OpenStudy (anonymous):

How do you simplify √(8x - 4 )?

ganeshie8 (ganeshie8):

looks you can factor 4 and pull it out of the radical

OpenStudy (anonymous):

so I would substitute -1 x 4?

ganeshie8 (ganeshie8):

not sure what you mean, can u do it and show me what you get

OpenStudy (anonymous):

Ok forget the radical for a second and factor this: 8x-4

OpenStudy (anonymous):

√(8x - 4) √(8x -(1x4) ?

ganeshie8 (ganeshie8):

looks good, keep going

OpenStudy (anonymous):

eek no: take out the 4- 4(2x-1) So now put that under the radical and you can take out the 4 right?

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

√(8x - 4) √(8x -(1x4) 2√(8x)?

OpenStudy (anonymous):

No, you don't ever get that. You get 8x-4 = 4(2x-1)

OpenStudy (anonymous):

And then you get 2 * sqrt(2x-1)

OpenStudy (anonymous):

i understand I think

OpenStudy (anonymous):

\[2 * \sqrt{2x-1}\] ok then that is basically the farthest you can simplify, have i helped?

OpenStudy (anonymous):

unless you have answer choices that say otherwise

OpenStudy (anonymous):

yes, but what if I have a larger problem like √11x - 6 ? how would the factoring work then?

TheSmartOne (thesmartone):

^ For that type of question, there would be no factoring to it... If they gave that equal to something, then we could solve for x, but factoring is a no.